# Modified Lepage-type test statistics for the weak null hypothesis

### v1

@@ Abstract
Detecting simultaneous shifts in the location and scale of two populations is a challenging problem in statistical research. A common way to address this issue is by combining location and scale test statistics. A well-known example is the Lepage test, which combines the Wilcoxon-Mann-Whitney test for location with the Ansari-Bradley test for scale. However, the Wilcoxon-Mann-Whitney test assumes that the population variances are equal, while the Ansari-Bradley test assumes the population medians are equal. This study introduces new approaches that combine recent methodological advances to relax these assumptions. We incorporate the Fligner-Policello test, a distribution-free alternative to the Wilcoxon-Mann-Whitney test that does not require the assumption of equal variances. The Fligner-Policello test is further enhanced by the Fong-Huang method, which provides an improved variance estimation. Additionally, we propose a new variance estimator for the Ansari-Bradley test, thereby eliminating the need for the equal medians assumption. These methodological modifications are integrated into the Lepage framework to operate under a weak null hypothesis. Simulation results suggest that these new tests are promising candidates for location-scale testing. The practical utility of the proposed tests is then demonstrated through an analysis of four real-world biomedical datasets. These empirical applications confirm the robustness and reliability of the modified tests for the two-sample independent location-scale problem.
Keywords: Fligner-Policello estimator; Fong-Huang estimator; Lepage-type statistics; Location-scale model; power of test; variance estimators. 11footnotetext: corresponding author: abid0100@gmail.com
@@ 1 Introduction
The comparison of treatment and control groups is a common practice in biomedical research and related disciplines. Standard two-sample tests, such as the Student’s tt–test and FF–ratio test, are frequently employed to evaluate the effects of different treatments, such as varying medication doses, on two independent groups. When the populations are normally distributed, these tests are considered the uniformly most powerful unbiased tests. However, the underlying assumption of normality for these tests is often not met in practice. Consequently, nonparametric two-sample tests are recommended, particularly in situations involving small sample sizes, where the verification of distributional assumptions is challenging or infeasible. These tests are preferred due to their lack of distributional assumptions within each sample and their demonstrated robustness under diverse conditions, as discussed by Blair and Higgins (1980) and Van der Vaart (2000). Among the numerous available statistical tests for comparing the locations of two independent samples, the nonparametric Wilcoxon-Mann-Whitney UU–test is frequently applied in a variety of settings (Letshedi et al. 2021; Lin et al. 2021 and Dao 2022). Ansari-Bradley test is a nonparametric test that is commonly used to compare the dispersion of two independent samples. It has become one of the core statistically analytical tools in most research and practical applications (Mukherjee and Marozzi 2019; Murakami and Neuhäuser 2025 and Hussain 2025).
Simultaneous variation in location and scale is a common occurrence in statistical analysis (see e.g., Neuhäuser 2002 and Murakami and Neuhäuser 2025). For instance, in randomized controlled trials, the null hypothesis of no treatment effect often assumes homogeneous variances among the different treatment groups. However, the presence of heteroscedasticity can itself be an indicator of an underlying treatment effect (Neuhäuser, 2002). A compelling example can be found in a clinical trial analyzed by Neuhäuser (2001), which investigated the frequency of exacerbations in patients with chronic obstructive pulmonary disease (COPD). In this context, a treatment effect could manifest a shift in both the location and scale parameters. This observation has spurred significant research into developing distribution-free tests capable of simultaneously detecting differences in both location and scale between two populations. A prominent example of such a test is the sum of the squared standardized Wilcoxon-Mann-Whitney UU–test and Ansari-Bradley CC–test statistics, a method introduced by Lepage (1971).
Statistical testing typically focuses on the strong null hypothesis, which posits an exact equality between populations. However, a weak null hypothesis, which concerns a relative marginal effect, can be relevant in certain situations (Brunner et al. 2002 and Fong and Huang 2019). The Wilcoxon-Mann-Whitney UU–test, for instance, cannot directly test this weak null hypothesis because the sampling distribution of its test statistic is not distribution-free under this condition (Chung and Romano 2013 and Chung and Romano 2016). To address this limitation, Fligner and Policello (1981) developed an alternative to the UU–test. They introduced a data-based approximation for the variance of the ranks. Their resulting test, known as the Fligner-Policello (FP) test, is calculated using the statistic $(U-1/2)/\sqrt{\widehat{Var}_{FP}(U)}$. Here, $\widehat{Var}_{FP}(U)$ represents a consistent variance estimator, which will be described in detail in Section 2. This test statistic can be approximated by a standard normal distribution or by a permutation distribution. When an additional distributional symmetry condition is met for both random variables $X$ and $Y$, the weak null hypothesis becomes equivalent to testing for the equality of population medians. Furthermore, the FP test is generally preferred over the UU–test when testing the weak null hypothesis without the assumption of homogeneous variances (e.g., Siegel and Castellan 1988; Zumbo and Coulombe 1997 and Mickelson 2013). The FP test, therefore, provides a robust alternative for comparing population locations under less restrictive assumptions.
In this paper, we propose several modifications to the Lepage test statistic to improve its power for the two-sample location-scale problem. These changes are specifically designed to maintain or increase statistical power under weak null hypotheses-that is, when unequal variances exist in location-testing or unequal medians exist in scale-testing. One of our approaches introduces the FP test as an alternative to the traditional Wilcoxon-Mann-Whitney for assessing the location component of the Lepage test. Further, we incorporate two key methodological advancements in variance estimation under less restrictive assumptions. First, we utilize the Fong-Huang (FH) estimator, which corrects a known error in the original FP estimator (Fong and Huang, 2019). This FH test enhances the accuracy to detect the location-difference, allowing it to be reliably compared to standard normal approximation as $(U-1/2)/\sqrt{\widehat{Var}_{FH}(U)}$. Second, we propose a novel empirical variance estimator for the Ansari-Bradley $C$ statistic. It can be implemented through standard normal approximation as $\big(C-E_{0}(C)\big)/\sqrt{\widehat{Var}(C)}$, eliminates the need for the assumption of equal medians. The Lepage-type test statistics developed in this study are constructed by incorporating the variance estimators of the $U$ and $C$ statistics. The subsequent sections of this paper provide the specific details and derivations of these new test statistics.
This study is structured as follows: In Section 2, we begin with the details of the $U$ statistic, its variance, and providing an overview of existing estimators proposed by Fligner and Policello (1981) and Fong and Huang (2019). Section 3 introduces the $C$ statistic, its variance, and a newly proposed variance estimator. Section 4 then presents the location-scale model and the Lepage-type statistics that form the basis of our analysis. Section 5 employs a Monte Carlo simulation to evaluate the performance of our proposed tests with classical Lepage test. In Section 6, we demonstrate the practical application of these tests using four real-world datasets. Finally, Section 7 summarizes the key findings of the research along with some future directions.
@@ 2.2 Variance of $U$ statistic
where $S_{Y}(X_{i})=E_{Y|X_{i}}\big(I(X_{i}<Y)\big)$. Thus we have
@@ 2.3 Fligner and Policello estimator
which is not equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$. However, this difference is negligible, being of order $O\big(m^{-1}n^{-1}\big)$.
@@ 2.4 Fong and Huang estimator
Fong and Huang (2019) recently proposed a corrected estimator for the variance of the $U$ statistic (3), defined as
@@
which is equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$, when $m\neq n$.
@@ 3.1 The classical $C$ statistic
The $C$ statistic, proposed by Ansari and Bradley (1960), is used to assess a two-sample dispersion problem. Its application requires the two populations to have identical medians. To calculate the statistic, the combined sample of $X$ and $Y$ is first ordered from least to greatest, represented as $\{Z_{1},~Z_{2},~\cdots,~Z_{N}\}$. Ranks are then assigned symmetrically from the tails of the ordered sample. Rank 1 is assigned to the minimum ($Z_{1}$) and maximum ($Z_{N}$) values, rank 2 to the second smallest ($Z_{2}$) and second largest ($Z_{N-1}$), and so on. The array of assigned ranks, denoted as $S_{N}$, depends on whether the combined sample size $N$ is even or odd. If $N$ is an even integer, the ranks are given by
$S_{N}=\bigg\{1,~2,~3,~\cdots,~\frac{N}{2},~\frac{N}{2},~\cdots,~3,~2,~1\bigg\}.$
@@
$S_{N}=\bigg\{1,~2,~3,~\cdots,\frac{N-1}{2},~\frac{N+1}{2},~\frac{N-1}{2},~\cdots,~3,~2,~1\bigg\}.$
@@
$\lambda_{i}=\begin{cases}1;&\ \ \text{if $Z_{i}\in$ $Y$ }\\ &\\ 0;&\ \ \text{otherwise,}\end{cases}$
@@ 3.2 Variance of $C$ statistic
$E_{0}\bigg(\frac{C}{n}\bigg)$ $=$ $\mu=\frac{2}{N}\sum_{i=1}^{N/2}i=\frac{\frac{2}{N}\times\frac{N}{2}\bigg(\frac{N}{2}+1\bigg)}{2}=\frac{N+2}{4}$
@@
These results are consistent with standard theory for the mean of a sample from a finite population (see, for example, Hollander et al. 2014).
For any NN (even or odd), the standardized statistic, $C^{*}$, is asymptotically normally distributed under $H_{0}$
$C^{*}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
Asymptotic normality results for $C^{*}$ are also available under general alternatives to $H_{0}$ (see, for example, Ansari and Bradley 1960 and Randles and Wolfe 1979).
@@ 3.3 A suggested variance estimator
This section presents a procedure for testing the dispersion of two independent samples, particularly when the underlying assumption regarding populations medians is violate. A key component of this method involves estimating the population variance, $\sigma^{2}$, using results from a sample. The expected value of this estimator is given by
$E(\hat{\sigma}^{2})=\sigma^{2}\bigg(1-\frac{1}{n}\times\frac{N-n}{N-1}\bigg)=\sigma^{2}\bigg(\frac{N(n-1)}{n(N-1)}\bigg).$
@@
$\widehat{Var}(C)=\hat{\sigma}^{2}\times\bigg(\frac{n(N-1)}{N(n-1)}\bigg)\times\bigg(\frac{n(N-n)}{N-1}\bigg)=\hat{\sigma}^{2}\times\bigg(\frac{n^{2}(N-n)}{N(n-1)}\bigg),$ (8)
where $\hat{\sigma}^{2}$ represents the observed variance of the ranks of sample $Y$. The proposed estimator is equally applicable for two different variances, as described in Eq. (6) and Eq. (3.2), regardless of whether the sample size $N$ is even or odd.
Following the strategy from Fong and Huang (2019), a simulation study is conducted to evaluate the performance of the proposed variance estimator. Data are generated from two different distributions
  1. $\bullet$
    A logistic distribution: $X\sim\mathcal{L}(0,1)$ versus $Y\sim\mathcal{L}(0,1)$,
  2. $\bullet$
    An exponential distribution: $X\sim\mathcal{E}(1)$ versus $Y\sim\mathcal{E}(1)$.
To compute the true variance of the $C$ statistic, Eq. (6) or (3.2) is used, depending on whether the sample size ($N$) is even or odd. This true variance is then compared to the average of the variance estimates obtained using Eq. (8). Table 1, based on 10,000 replications, shows that the proposed estimator provides accurate and consistent estimates of the true variance for both probability distributions. This consistency suggests that the proposed estimator is a consistent and distribution-free variance estimator of $C$ statistic. The test statistic for the weak null hypothesis is therefore given by
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}.$
Table 1: Comparison of $C$ statistic variance and its estimator.

### v3

# Enhanced Lepage-type test statistics for location-scale shifts with right-skewed data

@@ Abstract
Detecting simultaneous shifts in location and scale between two populations is a common challenge in statistical inference, particularly in fields like biomedicine where right-skewed data distributions are prevalent. The classical Lepage test, which combines the Wilcoxon-Mann-Whitney and Ansari-Bradley tests, can be suboptimal under these conditions due to its restrictive assumptions of equal variances and medians. This study systematically evaluates enhanced Lepage-type test statistics that incorporate modern robust components for improved performance with right-skewed data. We combine the Fligner-Policello test and Fong-Huang variance estimator for the location component with a novel empirical variance estimator for the Ansari-Bradley scale component, relaxing assumptions of equal variances and medians. Extensive Monte Carlo simulations across exponential, gamma, chi-square, lognormal, and Weibull distributions demonstrate that tests incorporating both robust components achieve power improvements of 10-25% over the classical Lepage test while maintaining reasonable Type I error control. The practical utility is demonstrated through analyses of four real-world biomedical datasets, where the tests successfully detect significant location-scale shifts. We provide practical guidance for test selection and discuss implementation considerations, making these methods accessible for practitioners in biomedical research and other disciplines where right-skewed data are common.
Keywords: Fligner-Policello test; Fong-Huang estimator; Lepage-type statistics; location-scale model; robust inference; right-skewed data; variance estimation.
@@ 1 Introduction
The comparison of treatment and control groups is a cornerstone of biomedical, psychological, and public health research. Practitioners often rely on standard two-sample tests, such as Student’s $t$-test for location and the $F$-ratio test for scale, to evaluate the effects of interventions. When data adhere to the assumption of normality, these parametric tests are among the most powerful available. However, this assumption is frequently violated in practice, particularly with small sample sizes where distributional properties are difficult to verify. In such scenarios, nonparametric tests are the recommended alternative due to their robustness and minimal distributional requirements (Mukherjee and Marozzi, 2019; Kumar et al., 2024).
Among nonparametric procedures, the Wilcoxon-Mann-Whitney (WMW) $U$-test is widely used for comparing the locations of two independent samples (Lin et al., 2021), while the Ansari-Bradley test is a common choice for assessing differences in scale (Mukherjee and Marozzi, 2019; Murakami and Neuhäuser, 2025; Hussain, 2025). A significant methodological challenge arises when a treatment effect manifests not as a pure location or scale shift, but as a combination of both—a scenario frequently encountered in practice (Murakami and Neuhäuser, 2025). For instance, in a clinical trial for chronic obstructive pulmonary disease, the therapeutic effect is characterized by simultaneous changes in both location and scale parameters (Neuhäuser, 2001). To address this, omnibus tests sensitive to any departure from the null hypothesis of identical distributions have been developed. A prominent distribution-free solution is the Lepage test (Lepage, 1971), which combines the standardized WMW $U$-test and Ansari-Bradley $C$-test statistics into a single quadratic form.
A critical, yet often overlooked, aspect of this problem is the prevalence of right-skewed data in many scientific fields. Outcome variables such as reaction times, biomarker concentrations, survival duration, and cost data frequently exhibit positive skewness (Fagerland, 2009). This distributional characteristic poses a particular challenge for conventional tests, whose power can be substantially diminished when the underlying populations are skewed and heteroscedastic. The standard Lepage test, while distribution-free, relies on the strong null hypothesis of identical population distributions (Lepage, 1971). This assumption is often untenable in the presence of heteroscedasticity, which itself may be an indicator of a meaningful treatment effect.
This limitation has motivated the development of tests for the weak null hypothesis, which concerns the equality of a relative effect size rather than the equality of the entire distribution (Fong and Huang, 2019). For instance, the (WMW) $U$-test is not directly applicable for testing this weak null, as the sampling distribution of its test statistic ceases to be distribution-free under this condition (Chung and Romano, 2013, 2016). In response to this limitation, Fligner and Policello (1981) developed an alternative to the $U$-test that relies on fewer assumptions concerning the underlying distributional forms of the populations. A parallel issue compromises the Ansari-Bradley test for scale differences, as its validity is contingent upon the assumption of equal medians—a condition that is often untenable and frequently incompatible with the location-scale alternatives under investigation. Modern robust methods (Kössler and Mukherjee, 2020) and permutation-based approaches (Chung and Romano, 2016) offer additional avenues for addressing heteroscedasticity and distributional asymmetry. However, a systematic evaluation of how these modern components perform when integrated into the Lepage framework—particularly for right-skewed data—remains lacking in the literature.
This study provides a comprehensive evaluation of enhanced Lepage-type test statistics that combine modern robust procedures with the classical Lepage framework. Rather than proposing fundamentally new methodology, our contribution lies in the systematic combination and evaluation of existing robust components for the specific challenge of right-skewed data. We integrate the Fligner-Policello test (Fligner and Policello, 1981) and its refined Fong-Huang variance estimator (Fong and Huang, 2019) for the location component, alongside a novel empirical variance estimator for the Ansari-Bradley statistic that relaxes the assumption of equal medians. By embedding these robustified components into the Lepage framework, we construct tests that are valid under weaker assumptions and exhibit enhanced sensitivity to the distributional shapes frequently encountered in practical applications.
The primary objectives of this study are threefold: first, to systematically evaluate the performance of various combinations of robust components within the Lepage framework; second, to provide practical guidance for test selection with right-skewed data across different sample size scenarios; and third, to demonstrate the practical utility of these methods through both extensive simulations and real biomedical applications. Our focus remains on right-skewed distributions commonly encountered in biomedical research, while acknowledging the need for further investigation with other distributional shapes.
The rest of this paper is organized as follows: Section 2 details the $U$ statistic and its robust variance estimators. Section 3 introduces the $C$ statistic and our proposed variance estimator, including theoretical justification. Section 4 presents the location-scale model and the Lepage-type test statistics. Section 5 presents an extensive Monte Carlo simulation study evaluating the performance of the tests across various distributional scenarios. Section 6 demonstrates the practical utility of the tests using real biomedical datasets, and Section 7 provides practical guidance, discusses limitations, and offers concluding remarks.
@@ 2.2 Variance of $U$ statistic
where $S_{Y}(X_{i})=E_{Y|X_{i}}\big(I(X_{i}<Y)\big)$. Thus, we have
@@ 2.3 Fligner and Policello estimator
which is not equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$. However, this difference is negligible, being of order $O\big(m^{-1}n^{-1}\big)$. The resulting test, known as the Fligner-Policello (FP) test, is based on the statistic $(U-1/2)/\sqrt{\widehat{Var}_{FP}(U)}$, which asymptotically follows a standard normal distribution (Hollander et al., 2014; Fong and Huang, 2019).
@@ 2.4 Fong and Huang estimator
Fong and Huang (2019) recently proposed a corrected estimator for the variance of the $U$ statistic (Eq. 3), defined as
@@
which is equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$, when $m\neq n$. The resulting Fong and Huang (FH) test, based on the standardized statistic $(U-1/2)/\sqrt{\widehat{Var}_{FH}(U)}$, provides a more accurate test for location differences via a standard normal approximation (Fong and Huang, 2019).
@@ 3.1 The classical $C$ statistic
The $C$ statistic, proposed by Ansari and Bradley (1960), is used to assess a two-sample dispersion problem. Its application requires the two populations to have identical medians. To calculate the statistic, the combined sample of $X$ and $Y$ is first ordered from least to greatest, represented as $\{Z_{1},\ Z_{2},\ \cdots,\ Z_{N}\}$. Ranks are then assigned symmetrically from the tails of the ordered sample. Rank 1 is assigned to the minimum ($Z_{1}$) and maximum ($Z_{N}$) values, rank 2 to the second smallest ($Z_{2}$) and second largest ($Z_{N-1}$), and so on. The array of assigned ranks, denoted as $S_{N}$, depends on whether the combined sample size $N$ is even or odd. If $N$ is an even integer, the ranks are given by
$S_{N}=\bigg\{1,\ 2,\ 3,\ \cdots,\ \frac{N}{2},\ \frac{N}{2},\ \cdots,\ 3,\ 2,\ 1\bigg\}.$
@@
$S_{N}=\bigg\{1,\ 2,\ 3,\ \cdots,\frac{N-1}{2},\ \frac{N+1}{2},\ \frac{N-1}{2},\ \cdots,\ 3,\ 2,\ 1\bigg\}.$
@@
$\lambda_{i}=\begin{cases}1;&\text{if $Z_{i}\in Y$}\\ \\ 0;&\text{otherwise},\end{cases}$
@@ 3.2 Variance of $C$ statistic
$E_{0}\biggl(\frac{C}{n}\biggr)$ $=\mu=\frac{2}{N}\sum_{i=1}^{N/2}i=\frac{\frac{2}{N}\times\frac{N}{2}\biggl(\frac{N}{2}+1\biggr)}{2}=\frac{N+2}{4}$
@@
These results are consistent with standard theory for the mean of a sample from a finite population (Hollander et al., 2014). For any $N$ (even or odd), the standardized statistic, $C^{*}$, is asymptotically normally distributed under $H_{0}$
$C^{*}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\xrightarrow{d}\mathcal{N}(0,1).$
Asymptotic normality results for $C^{*}$ are also available under general alternatives to $H_{0}$ (Ansari and Bradley, 1960; Randles and Wolfe, 1979).
@@ 3.3 A suggested variance estimator
To relax the assumption of equal medians, we propose an empirical variance estimator for $C$. Let $\hat{\sigma}^{2}$ be the observed variance of the ranks assigned to the $Y$ sample. The expected value of this estimator is given by
$E(\hat{\sigma}^{2})=\sigma^{2}\bigg(1-\frac{1}{n}\cdot\frac{N-n}{N-1}\bigg)=\sigma^{2}\bigg(\frac{N(n-1)}{n(N-1)}\bigg).$
@@
$\widehat{Var}(C)=\hat{\sigma}^{2}\cdot\bigg(\frac{n(N-1)}{N(n-1)}\bigg)\cdot\bigg(\frac{n(N-n)}{N-1}\bigg)=\hat{\sigma}^{2}\cdot\frac{n^{2}(N-n)}{N(n-1)}.$ (8)
This estimator is equally applicable for two different variances, as described in Eq. (6) and Eq. (3.2), regardless of whether the sample size $N$ is even or odd.
3.4 Theoretical properties of the proposed estimator
Theorem 1 (Consistency of $\widehat{Var}(C)$)
Under the weak null hypothesis and assuming continuous distributions with finite fourth moments, the proposed estimator $\widehat{Var}(C)$ is consistent for $Var_{0}(C)$.
Proof
Let $\hat{\sigma}^{2}$ be the sample variance of the ranks assigned to the $Y$ sample. Under the weak null hypothesis, $\hat{\sigma}^{2}$ is a consistent estimator of $\sigma^{2}$ by the weak law of large numbers. The finite population correction terms $\frac{n(N-1)}{N(n-1)}$ and $\frac{n(N-n)}{N-1}$ are deterministic functions of $m$ and $n$ that converge to 1 as $m,n\to\infty$ with $m/N\to\gamma\in(0,1)$. By the continuous mapping theorem, the product $\widehat{Var}(C)$ converges in probability to $\sigma^{2}$, which equals $Var_{0}(C)$ under the null. $\square$
Proposition 1 (Finite-sample bias)
The proposed estimator $\widehat{Var}(C)$ is unbiased for $Var_{0}(C)$ under the strong null hypothesis.
Proof
Under the null hypothesis, we have
$E[\hat{\sigma}^{2}]=\sigma^{2}\left(\frac{N(n-1)}{n(N-1)}\right).$
Substituting into the estimator (8)
$E[\widehat{Var}(C)]$ $=E[\hat{\sigma}^{2}]\cdot\frac{n^{2}(N-n)}{N(n-1)}$
$=\sigma^{2}\left(\frac{N(n-1)}{n(N-1)}\right)\cdot\frac{n^{2}(N-n)}{N(n-1)}$
$=\sigma^{2}\cdot\frac{n(N-n)}{N-1}=Var_{0}(C).$
$\square$
3.5 Empirical validation
Following the strategy from Fong and Huang (2019), we conducted a simulation to evaluate the performance of $\widehat{Var}(C)$ under the null hypothesis for logistic and exponential distributions. Table 1 shows that the average of our estimates over 10,000 replications is virtually identical to the true analytical variance, $Var_{0}(C)$, across various sample sizes. This demonstrated accuracy, coupled with the estimator’s theoretical properties, confirms that $\widehat{Var}(C)$ serves as a robust distribution-free variance estimator for the $C$ statistic.
3.6 Construction of the test statistic
Given the accuracy of $\widehat{Var}(C)$, we propose the following standardized statistic for testing the weak null hypothesis
$C^{*}_{P}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}.$
A permutation study confirmed that the critical values for $C^{*}_{P}$ and the classical $C^{*}$ are numerically identical for various sample sizes, supporting the use of $C^{*}_{P}$. Specifically, the critical values are found at 5% upper-tailed to be 1.697056 for $m=n=5$, 1.669331 for $m=n=10$, 1.674574 for $m=20,n=10$, and 1.62453 for $m=n=20$.
3.7 Asymptotic normality via Slutsky’s Theorem
The asymptotic normality of $C^{*}_{P}$ under the weak null hypothesis follows from Slutsky’s Theorem (see e.g., Casella and Berger 2002, p. 239–240). We can write
$C^{*}_{P}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\cdot\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}.$ (9)
The first term, $\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}$, converges in distribution to a standard normal variable by the classical asymptotic theory of the Ansari-Bradley test. The second term, $\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}$, converges in probability to 1 due to the consistency of $\widehat{Var}(C)$ established in Theorem 1. Therefore, by Slutsky’s Theorem, $C^{*}_{P}\xrightarrow{d}\mathcal{N}(0,1)$.
Table 1: Comparison of the true variance of the $C$ statistic and the proposed estimator.